OpenStudy (anonymous):

olve 4x + 10y = 94 y = 3x – 28

OpenStudy (anonymous):

*solve

jimthompson5910 (jim_thompson5910):

Start with 4x + 10y = 94 then replace y with 3x-28 to go from 4x + 10y = 94 to 4x + 10(3x-28) = 94

jimthompson5910 (jim_thompson5910):

Then solve 4x + 10(3x-28) = 94 for x

OpenStudy (anonymous):

x=11, y=5